OpenStudy (ledah):

solve using substitution 3x + 2y = -4 x + 2y = -8

jhonyy9 (jhonyy9):

x=-8-2y so this you need to substitute in first equation in place of x

OpenStudy (anonymous):

\[x+2y =-8\]\[x=-8+2y\]\[3(-8+2y)+2y=-4\]\[-24+6y+2y=-4\]\[-24+8y=-4\]\[8y=20 \rightarrow y=\frac{20}{8}=2.5\]\[x+2(2.5)=-8\]\[x=-8-5 \rightarrow x= -13\] now check the result by substituting back into the first equation\[-13+(2 \times 2.5)=-8\]\[-13+5=-8\] and it does